OpenStudy (anonymous):

what is standard form? and how do I sovle through: (-4,3), parallel to y=-5/4x-3

hartnn (hartnn):

parallel lines have same slope so what is the slope of the required line ?

OpenStudy (anonymous):

so they have the same slope?

OpenStudy (anonymous):

then to perpendicular have recpical slope?

hartnn (hartnn):

negative reciprocal to be precise. -1/(slope)

OpenStudy (anonymous):

okay but how do I solve through: (-4,3), parallel to y=-5/4x-3

OpenStudy (anonymous):

in standard form

hartnn (hartnn):

whats the slope of line y=-5/4x-3 ? u know how to find ?

OpenStudy (anonymous):

-5/4

hartnn (hartnn):

thats correct! so , since parallel lines have same slope , u got the slope of your line as -5/4 now, put that in any of the line equation form you know , then convert it into standard form

hartnn (hartnn):

standard form is ax+by+c=0

OpenStudy (anonymous):

so it would be like -4x+3y+-5/4=0

OpenStudy (anonymous):

Im lost

hartnn (hartnn):

which form of line u know ? u know y=mx+b ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i need to have it in standard form

OpenStudy (anonymous):

in it y=-5/4x-2

hartnn (hartnn):

m=slope = -5/4 now put the point in that to find b. we'll convert it to standard form after u find b

OpenStudy (anonymous):

in b=-2

OpenStudy (anonymous):

or -5

hartnn (hartnn):

yes, thats correct. y=-5/4x-2 now we need to bring that in form of ax+by+c=0 multiply 4 on both sides to eliminate fraction

OpenStudy (anonymous):

no wait i mean it's b=-8

hartnn (hartnn):

b= -2 was correct. 3= -5/4(-4)+b gives b=-2

OpenStudy (anonymous):

oh okay

hartnn (hartnn):

y=-5/4x-2 now we need to bring that in form of ax+by+c=0 multiply 4 on both sides to eliminate fraction

OpenStudy (anonymous):

oh so it would be 4y=-5x-2

hartnn (hartnn):

-2*4 = -8

OpenStudy (anonymous):

oh so do I always have to multipy it to the b

hartnn (hartnn):

4 was to be multiplied to every term , both sides.

OpenStudy (anonymous):

oh okay then what is the next step

hartnn (hartnn):

u get 4y= -5x-8, right ? add 5x+8 to both sides to get 0 on right side

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what do I do what do u mean add it

hartnn (hartnn):

4y+5x+8 = -5x-8+5x+8 = 0 so u get 5x+4y + 8= 0 <---standard form

OpenStudy (anonymous):

but I thought the answer was 5x+4y=-8

hartnn (hartnn):

yes! thats is standard form . ax+by =c so you are correct.

OpenStudy (anonymous):

oh okay thank you so much and sorry for the hard time

hartnn (hartnn):

no problem. welcome ^_^